Construct the truth tables for the following compound propositions.




1. (π‘Ž ∧ 𝑏) β†’ (π‘Ž ∨ 𝑏)




2. (π‘Ž β†’ 𝑏) ∨ (∼ π‘Ž β†’ 𝑏)




3. [(π‘Ž β†’ 𝑏) ∧ (π‘Ž β†’ 𝑐)] β†’ (π‘Ž β†’ (𝑏 ∧ 𝑐))




4. ∼ (𝑏 ∧ π‘Ž) ↔ (𝑏 ∨∼ π‘Ž)




5. (π‘Ž β†’ 𝑏) ∨ 𝑐

1
Expert's answer
2022-03-07T17:18:03-0500

1.

"\begin{array}{|c|c|c|c|c|}
\hline a & b & a \wedge b & a \vee b & (a \wedge b) \rightarrow(a \vee b) \\
\hline \text { T } & \text { T } & \text { T } & \text { T } & \text { T } \\
\hline \text { T } & \text { F } & \text { F } & \text { T } & \text { T } \\
\hline \text { F } & \text { T } & \text { F } & \text { T } & \text { T } \\
\hline \text { F } & \text { F } & \text { F } & \text { F } & \text { T } \\
\hline
\end{array}"


2.

"\begin{array}{|c|c|c|c|c|c|}
\hline a & b & a \rightarrow b & \neg a & \neg a \rightarrow b & (a \rightarrow b) \vee(\neg a \rightarrow b) \\
\hline \text { T } & \text { T } & \text { T } & \text { F } & \text { T } & \text { T } \\
\hline \text { T } & \text { F } & \text { F } & \text { F } & \text { T } & \text { T } \\
\hline \text { F } & \text { T } & \text { T } & \text { T } & \text { T } & \text { T } \\
\hline \text { F } & \text { F } & \text { T } & \text { T } & \text { F } & \text { T } \\
\hline
\end{array}"


3.

"\begin{array}{|c|c|c|c|c|c|c|c|c|}
\hline a & b & c & a \rightarrow b& a \rightarrow c & b \wedge c & a \rightarrow & (a \rightarrow b) \wedge & ((a \rightarrow b) \wedge(a \rightarrow c)) \rightarrow \\
& & & & & & (b \wedge c) & (a \rightarrow c) & (a \rightarrow(b \wedge c)) \\
\hline \text { T } & \text { T } & \text { T } & \text { T } & \text { T } & \text { T } & \text { T } & \text { T } & \text { T } \\
\hline \text { T } & \text { T } & \text { F } & \text { T } & \text { F } & \text { F } & \text { F } & \text { F } & \text { T } \\
\hline \text { T } & \text { F } & \text { T } & \text { F } & \text { T } & \text { F } & \text { F } & \text { F } & \text { T } \\
\hline \text { T } & \text { F } & \text { F } & \text { F } & \text { F } & \text { F } & \text { F } & \text { F } & \text { T } \\
\hline \text { F } & \text { T } & \text { T } & \text { T } & \text { T } & \text { T } & \text { T } & \text { T } & \text { T } \\
\hline \text { F } & \text { T } & \text { F } & \text { T } & \text { T } & \text { F } & \text { T } & \text { T } & \text { T } \\
\hline \text { F } & \text { F } & \text { T } & \text { T } & \text { T } & \text { F } & \text { T } & \text { T } & \text { T } \\
\hline \text { F } & \text { F } & \text { F } & \text { T } & \text { T } & \text { F } & \text { T } & \text { T } & \text { T } \\
\hline
\end{array}"


4.

"\begin{array}{|c|c|c|c|c|c|c|}
\hline b & a & b \wedge a & \neg a & \neg(b \wedge a) & b \vee \neg a & \neg(b \wedge a) \leftrightarrow(b \vee \neg a) \\
\hline \text { T } & \text { T } & \text { T } & \text { F } & \text { F } & \text { T } & \text { F } \\
\hline \text { T } & \text { F } & \text { F } & \text { T } & \text { T } & \text { T } & \text { T } \\
\hline \text { F } & \text { T } & \text { F } & \text { F } & \text { T } & \text { F } & \text { F } \\
\hline \text { F } & \text { F } & \text { F } & \text { T } & \text { T } & \text { T } & \text { T } \\
\hline
\end{array}"


5.

"\begin{array}{|c|c|c|c|c|}
\hline a & b & c & a \rightarrow b & (a \rightarrow b) \vee c \\
\hline \text { T } & \text { T } & \text { T } & \text { T } & \text { T } \\
\hline \text { T } & \text { T } & \text { F } & \text { T } & \text { T } \\
\hline \text { T } & \text { F } & \text { T } & \text { F } & \text { T } \\
\hline \text { T } & \text { F } & \text { F } & \text { F } & \text { F } \\
\hline \text { F } & \text { T } & \text { T } & \text { T } & \text { T } \\
\hline \text { F } & \text { T } & \text { F } & \text { T } & \text { T } \\
\hline \text { F } & \text { F } & \text { T } & \text { T } & \text { T } \\
\hline \text { F } & \text { F } & \text { F } & \text { T } & \text { T } \\
\hline
\end{array}"